OpenStudy (anonymous):

is the imporer integral \[\int\limits\limits_{0}^{+\infty}re ^{-r ^{2}}dr\] convorgent? Justify your answer

OpenStudy (anonymous):

i'm trying to solve it but i am not soon to the answer yet -.-

OpenStudy (anonymous):

lets take u=r^2 here will make the thing easier

OpenStudy (anonymous):

now we have \[\int\limits_{0}^{+\infty}re^{-u}dr\]

OpenStudy (anonymous):

and now i guess we can do that : \[1/2 \int\limits_{0}^{+\infty}e ^{-u}du\]

OpenStudy (anonymous):

k now we got \[(-1/2)e^{-r ^{2}} + c\]

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

making u=-r^2 would have been a little bit easier

OpenStudy (anonymous):

\[\int\limits_{0}^{+\infty}re ^{-r ^{2}}dr = \lim_{l \rightarrow \infty}\int\limits_{0}^{1}re ^{-r ^{2}}dr\]

OpenStudy (anonymous):

as r goes to infinity 1/r^2 goes to 0

OpenStudy (anonymous):

now i think we can say \[\lim_{l \rightarrow \infty}1/2 [-e ^{-l ^{2}}]^l\]

OpenStudy (anonymous):

yeapp

OpenStudy (anonymous):

\[-\frac{1}{2}e^{\frac{1}{r^2}}\] \[-\frac{1}{2}e^0\] -1/2

OpenStudy (anonymous):

and now lets work on it.... ok now we get -1/2[0-2] = 1/2

OpenStudy (anonymous):

so the given integral converges to 1/2 i guess

OpenStudy (anonymous):

yes 1/2

OpenStudy (anonymous):

how old are you man? :D

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

no reason :D there shuld not everything has a reson

OpenStudy (anonymous):

19

OpenStudy (anonymous):

me = 14

OpenStudy (anonymous):

14 and already doing integral, nice!

OpenStudy (anonymous):

yeap :D 3 mounths ago, i did not kno even limits, :D it took only three mounths me to learn algebra 1-2, calculus 1-2 :D not seeen a really hard stuff yet :(

OpenStudy (anonymous):

doing calculus is not hard, applying it is somewhat difficult

OpenStudy (anonymous):

yes :D exatly i sometimes need calculus in my 2D games, and it really beings hard to use the things i know -.-' really i always need to have a look at books ...

OpenStudy (anonymous):

games, tell me about that

OpenStudy (anonymous):

well, i begin to work on sof. dev. when i was 12, learned functions when i was 12 :D well i begin with QBASIC, solved over 6.000 algorithm quesitons learned, C, C++, C#, VB.GWBASIC, some java, PHP ... nowadays i work on Assembly X86 :D , i developed over 4 o 5 games in C#,XNA , wrote all the algorithms that i need in my game on my own (for example to make a cannon rotate :D ) nowadays i am trying to learn C++ with DirectX too :D to develop better games :D but i need to work calculus 3 much thrre dimensional space to start to work on 3D game programming :D :D thats why i just stopped to work on game programming much and started to work on maths last three mounths :D and learned limit, differential , contiunity, integral ... etc thousends of things in these threee mounths in 2 or 1 more mounths ill be able to start to work on three dimensional space, :D and in the mean time i learn about DirectX so ill contiune at where i stopped on game programming :D

OpenStudy (anonymous):

and i think its the hardest question i ever solved : http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e6a40800b8b4a2b95d45a89